OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -14, and 5 + 8i

OpenStudy (tkhunny):

Well, get to work! f(x) = (x-4)(x+14)(x-(??))(x+(??)) You tell me the missing pieces!

OpenStudy (anonymous):

5+8i and 5-8i

OpenStudy (tkhunny):

Fooled you. That would be exactly correct if I had written it exactly correctly. I'll try again and fill in your CORRECT answers. f(x) = (x-4)(x+14)(x-(5+8i))(x-(5-8i)) That's all there is to it. So start multiplying to get the standard form.

OpenStudy (anonymous):

i dont remember how to multiply out the imaginary part

OpenStudy (tkhunny):

There's nothing to remember. It is just like any other multiplication. Do you know this? \((a+b)(a-b) = a^{2}-b^{2}\) That's what we're doing. (x-(5+8i))(x-(5-8i)) = (x-5-8i)(x-5+8i) Just multiply those trinomials. There's no secret. Just match up all the pieces.

OpenStudy (anonymous):

im confused

OpenStudy (tkhunny):

Can you multiply these two numbers? 123*456?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i know how to do that

OpenStudy (tkhunny):

How did you do it? 6*3 6*2 6*1 5*3 4*2 5*1 4*3 4*2 4*1 In an organized fashion? That is EXACTLY how to multiply the trinomial you are facing.